Using a 0.25 M phosphate buffer with a pH of 7.0, you add 0.79 mL of 0.49 M HCl to 42 mL of the buffer. What is the new pH of the solution?

Using a 0.25 M phosphate buffer with a pH of 7.0, you add 0.79 mL of 0.49 M NaOH to 42 mL of the buffer. What is the new pH of the solution?
Answer #1

The molar ratio of the buffer is calculated, using the Henry Hasselbach equation cleared:

n Salt / n Acid = 10 ^ (pH - pKa) = 10 ^ (7 - 7.21) = 0.62

It has:

i) n Salt - 0.62 * n Acid = 0

ii) n Salt + n Acid = M * V = 0.25 M * 0.042 L = 0.0105 mol

System of equations is applied and you have:

n Salt = 0.004 mol

n Acid = 0.006 mol

The added moles of HCl are calculated:

n HCl = 0.49 * 0.00079 = 0.0004 mol

HCl reacts with salt (decreasing it) and forms acid (increasing it):

The new pH is calculated:

pH = pKa + log (n Salt / n Acid) = 7.21 + log (0.004 - 0.0004 / 0.006 + 0.0004) = 6.77

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.